I need help with all these equations, Please!
Plot the graph of 3x + 15 = x + 25
Plot the graph of 4x – 3y = 10
Plot the graph of y4 + 3y2 -4 = 0
Plot the graph of 3x2 + 5 + 7x

Answer by stanbon(48564)

Plot the graph of 3x + 15 = x + 25
2x = 10
x = 5
Draw a vertical line thru all the points where x = 5
========================================================
Plot the graph of 4x – 3y = 10
Solve for y: y = (4/3)x - (10/3)
You need two points:
Let x = 0 then y = -(10/3) giving point (0,-10/3)
Let x = 3 then y = 4-10/3 = 2/3 giving point (3,2/3)
Plot the two points and draw a line thru them.
=========================================================
Plot the graph of y^4 + 3y^2 -4 = 0
Factor: (y^2+4)(y^2-1) = 0
True if y = 1 or -1
----------------------------------------
Plot the graph of 3x2 + 5 + 7x
Is this supposed to be y = 3x^2+7x+5 ?
If it is, the graph is a parabola that opens up (instead of down).
The bottom point (vertex) is at (-7/6,0.91)
========================================
Cheers,
stan H.